At equilibrium, the concentrations in this system were found to be[N


2

]=[O


2

]=0.200 M





and[NO]=0.500 M.



N


2

(g)+O


2

(g)↽−

−⇀

2NO(g)












If more NO

 is added, bringing its concentration to 0.800 M,

 what will the final concentration of NO

 be after equilibrium is re‑established?


1
Expert's answer
2022-03-08T15:25:03-0500

Keq = [NO]2/[N2]*[O2] = (0.5)2/(0.2*0.2) = 6.25


Let the concentration of N2 and O2 increase by x M after after equilibrium is re‑established. 


N2(g)  +  O2(g)  =  2NO(g)

0.2+x + 0.2+x = 0.8 - 2x


Keq = (0.8-2x)2/(0.2 + x)2 = 6.25


After solving this the equation we get x = 0.067 M


So final concentration of NO be after equilibrium is re‑established = 0.800 - 2*0.067 = 0.667 M